19. Hassan keeps picking playing cards out of a standard deck of 52 cards, hoping that he will (1 point) draw a queen. There are four queens in the deck. After looking at each card, he places it back in the deck. What is the probability that Hassan will draw a queen in the first seven attempts? 043% 046% 57% 68%
Answer #1

The probability of getting a queen in a randomly selected draw

52 13

Let X denotes the number of queen drawn in the first seven attempts.

X ~ Binomial(n = 7, p = 1/13)

The probability mass function of X is

*G) * (1-187-r , x = 0, 1,2,3, T r13

The probability that Hassan will draw a queen in the first seven attempts

= P(X > 1)

=1 - P(X = 0)

= 1 - (1 - rac{1}{13})^{7 - 0}

1-0.571038

0.428962

0.43

43%